# Landscape of Thoughts: Visualizing the Reasoning Process of Large Language Models

### v1

@@ 2.3 Metrics
Besides the qualitative 2D visualization, we introduce three quantitative metrics to help understand the behavior of the LLM at different reasoning steps, which are defined on the intermediate states introduced in Sec. 2.2.
@@
Consistency. To understand whether the LLM knows the answer before generating all thoughts, we compute the consistency of state $s_{i}$ by checking whether ${\bm{s}}_{i}$ and ${\bm{s}}_{n}$ agree
$\text{Consistency}(s_{i})=\text{1}(\mathop{\mathrm{argmin}}{\bm{s}}_{i}=\mathop{\mathrm{argmin}}{\bm{s}}_{n}).$ (2.6)
@@
Uncertainty. To know how confident the LLM is about its predictions at intermediate steps, we compute the uncertainty of state $s_{i}$ as the entropy of ${\bm{s}}_{i}$ (note $\sum_{d\in{\bm{s}}_{i}}d=1$)
$\text{Uncertainty}(s_{i})=-\sum_{d\in{\bm{s}}_{i}}d\cdot\log d.$ (2.7)
@@
$\text{Perplexity}(t_{i})=p_{\text{LLM}}(t_{i}|s_{i-1})^{-1/|t_{i}|}.$ (2.8)
@@ 3 Results and Observations
In this section, we utilize the landscape of thoughts to analyze the reasoning behavior of LLMs. Specifically, we conduct a comprehensive evaluation and extract several insightful observations by comparing the landscape of thoughts across (1) reasoning algorithms (Sec. 3.1), (2) reasoning tasks (Sec. 3.2), and (3) language models (Sec. 3.3).
To help understand the qualitative visualizations, we quantitatively calculate the consistency and uncertainty of states, as well as the perplexity of thoughts, all previously introduced in Sec. 2.3. Unless stated otherwise, we employ Llama-3.1-70B with CoT as the default configuration in evaluations. More visualization cases are in Appendix B.
3.1 Comparison across Reasoning Algorithms
Setup. We evaluate the default model with four reasoning algorithms: chain-of-thought (CoT) (Wei et al., 2022), least-to-most (LtM) (Zhou et al., 2023), MCTS (Zhang et al., 2024), and tree-of-thought (ToT) (Yao et al., 2023a). We run these algorithms on 50 problems randomly selected from the AQuA dataset. The corresponding landscapes are presented in Fig. LABEL:fig:understanding_diff_methods, which yield the following observations.
Observation 3.1 (The landscapes converge faster to the correct answers are of higher reasoning accuracy).
By comparing the four groups of landscapes in Fig. LABEL:fig:understanding_diff_methods, we observe that the states scatter dispersedly at early stages and gradually converge to correct (or incorrect) answers in later stages. Here, converge means the trend of a reasoning path approaching one answer. As can be seen from Fig. LABEL:fig:understanding_diff_methods, different reasoning algorithms present diverse landscapes. Generally, methods with more scattered landscapes (converge slower) present lower accuracy than those that converge faster.
Observation 3.2 (Wrong paths quickly converge to wrong answers, while correct paths slowly step to correct answers).
By comparing the landscapes of failure and success paths, it is found that the failure paths usually converge to the wrong answers at 20-40% states. By contrast, the states in the success paths converge to the correct answers at 80-100% states. This implies that early states can lead to any potential answers (from model perspective), while the correct answers are usually determined at the end of reasoning paths.
Observation 3.3 (Compared to failure paths, the intermediate states in correct paths have higher consistency w.r.t. the final state).
By comparing the consistency plots in Fig. LABEL:fig:understanding_diff_methods, we found that the model generally has low consistency between the intermediate states and the final state. Notably, the consistency of wrong paths is significantly lower than that of correct paths. This implies that the reasoning process can be quite unstable. Even though decoding algorithms like CoT and LtM are designed to solve a problem directly (without explorations), the generated thoughts by these methods do not consistently guide the reasoning path to the answer.
@@ 3.2 Comparison across Reasoning Tasks
Setup. Besides the AQuA, we include MMLU, CommonsenseQA, and StrategyQA datasets. We run the base model with CoT on 50 problems per dataset. The observations followed are derived from the landscapes in Fig. LABEL:fig:understanding_diff_datasets.
@@
The landscapes of AQuA, MMLU, and StrategyQA exhibit organized search behavior with higher state diversity, while CommonSenseQA presents concentrated search regions, reflecting direct knowledge retrieval rather than step-by-step reasoning processes. These distinct landscape patterns demonstrate the potential to reveal underlying domain relationships across different reasoning tasks.
@@
The histograms in Fig. LABEL:fig:understanding_diff_datasets show that path perplexity consistently increases during reasoning across all datasets. AQuA and MMLU show distinctly higher levels of uncertainty. For StrategyQA, correct paths show increasing consistency that surpasses incorrect paths at around 60% states, while incorrect paths show decreasing consistency. However, extending beyond the typical three-step requirement (Geva et al., 2021), the later stages (60-100% states) show increasing perplexity as well as lower uncertainty.
3.3 Comparison across Language Models
Setup. In this part, we study several LLMs’ behavior across different parameter scales (1B, 3B, 8B, and 70B). We run each model with CoT on 50 problems from the AQuA dataset. The landscapes of these models are shown in Fig. LABEL:fig:understanding_diff_models.
Observation 3.6 (The landscape converges faster as the model size increase).
As model parameters scale from 1B to 70B, the corresponding landscape demonstrates faster convergence to the correct answers with higher density in the last 20% states, aligning with the increasing accuracy. With more parameters to store information, larger models can access broader knowledge (Allen-Zhu and Li, 2024). This leads to more confident solutions, demonstrated by more focused answer patterns and lower uncertainty.
Observation 3.7 (Larger models have higher consistency, lower uncertainty, and lower perplexity).
As the model size increases, the consistency increases, at the same time, the uncertainty and perplexity decrease significantly. This also aligns with the higher accuracy for the large models.
@@ 4 Adapting Visualization to Predictive Models
One advantage of our method is that it can be adapted to a model to predict any property users observe. Here, we show how to convert our method to a lightweight verifier for voting reasoning paths, following the observations in Sec. 3. Note that this methodology is not limited to verifiers. Users can use the same technique to adapt the visualization tool to monitor other properties in their own scenarios.
@@ 4.1 A Lightweight Verifier
Observation 3.2 and 3.3 show that the convergence speed and consistency of intermediate states can distinguish correct and wrong paths. Inspired by these observations, we build a model $f:{\mathbb{R}}^{(k+1)\times n}\to\{0,1\}$ to predict the correctness of a reasoning path based on the state features $\{{\bm{s}}_{i}\}_{i=1}^{n}$ and consistency metric $\{\text{Consistency}(s_{i})\}_{i=1}^{n}$. The insight is that the state features, used to compute the 2-D visualization, encode rich location information of the states and can be used to estimate the convergence speed. Due to the small dimensionality of these features, we parameterize $f$ with a random forest (Breiman, 2001) to avoid overfitting.
We use this model as a verifier to enhance LLM reasoning (Cobbe et al., 2021). Unlike popular verifiers (Lightman et al., 2023) that involve a moderately sized language model on textual thoughts, our verifier operates on state features and is super lightweight. We train a verifier on thoughts sampled on the training split of each dataset and apply it to vote reasoning paths at test time. Given $q$ paths sampled by a decoding algorithm, the final prediction is produced by a weighted majority voting given as follows:
$\widehat{y}=$ $\mathop{\mathrm{argmax}}_{c\in{\mathcal{C}}}\sum_{i=1}^{q}$ $\text{1}(\widehat{y}^{(i)}=c)\cdot f(\{{\bm{s}}_{i}\}_{i=1}^{n},\{\text{Consistency}(s_{i})\}_{i=1}^{n}).$ (4.1)
Figure 5: The accuracy of reasoning (averaging across all four datasets). Results for each dataset are in Appendix B.4.
Figure 6: Demonstration of the inference-time scaling effect of the verifier. We show the voting accuracy (%) on StrategyQA scales with the number of reasoning paths.
@@ 4.2 Experimental Results
We evaluate our numerical verifier against an unweighted voting baseline (Wang et al., 2023b) with various language models, decoding algorithms, and reasoning datasets. Detailed experimental settings are in Appendix B.1.
@@
Effectiveness of the verifier. We first compare our verifier against the unweighted voting baseline, each applied to 10 reasoning paths. As shown in Fig. 5, our verifier consistently enhances the reasoning performance of all models and decoding algorithms, even though our verifier does not use any pre-trained language model. Notably, smaller language models (1B and 3B) show significant performance gains with the verifier’s assistance, achieving substantial improvements over their original capabilities of reasoning.
@@
Test-time scaling. While the improvement of the verifier seems marginal with 10 reasoning paths, our verifier can provide substantial performance gain with more reasoning paths. We adjust the number of reasoning paths from 1 to 50, and plot the results of the verifier and the unweighted voting baseline in Fig. 6. Models with our verifier exhibit significantly stronger scaling behaviors, achieving over 65% accuracy. In contrast, the performance of the baseline saturated around 30% accuracy. These results suggest that our state features, which are used in both the visualization tool and the verifier, capture important information about the reasoning behavior of LLMs. Thus, the verifier can boost test-time scaling, especially in solving complex problems.
@@
Cross-dataset and cross-model transferability. One interesting property of the state features and metrics is that their shape and range are agnostic to the model and dataset, suggesting that we may deploy the verifier trained on one dataset or model in another setting. As illustrated in Fig. LABEL:fig:_transfer, we evaluate how the verifier transfers across reasoning datasets (e.g., train on AQuA and test on MMLU) and model scales (e.g., train on 1B model and test on 70B model). We observe some positive transfers across datasets and models. For example, a verifier trained on AQuA can improve the performance of StrategyQA by 4.5%. A verifier trained on the 70B model also improves the performance of the 3B model by 5.5%. However, some cases do not benefit from the transferring verifiers. We leave improving the transferability of the state features and metrics as future work.
5 Related Work
Reasoning with large language models. Chain-of-Thought (CoT) prompting (Wei et al., 2022; Kojima et al., 2022) has empowered LLMs to tackle multi-step reasoning problems by generating intermediate steps before producing a final answer. Building upon CoT, numerous methods have been proposed to address various challenges, including compositional generalization (Zhou et al., 2023; Khot et al., 2023), planning (Yao et al., 2023a; Hao et al., 2023), and rule learning (Zhu et al., 2023) within the CoT reasoning.
Beyond solving reasoning tasks, CoT has also emerged as a foundational framework for other techniques, such as fine-tuning LLMs (Zelikman et al., 2022), enabling LLM-based agents (Yao et al., 2023b), and facilitating test-time scaling (Snell et al., 2024). Nevertheless, most of these approaches are developed in a trial-and-error manner, largely due to the absence of proper tools for analyzing the CoT.
Understanding chain-of-thought reasoning. There are a few studies that explore what makes CoT prompting effective by perturbing its exemplars. To be specific, Madaan and Yazdanbakhsh (2022) found that the text and patterns of exemplars help CoT generate sentences resembling correct answers. Besides, Wang et al. (2023a) highlighted the importance of maintaining the correct order of reasoning steps, while Ye et al. (2022) demonstrated that using complementary exemplars can enhance reasoning performance. Furthermore, CoT can benefit from longer reasoning chains, even without new information to the prompt (Jin et al., 2024).
Another line of research investigates CoT’s general behavior (Tang et al., 2023; Saparov and He, 2023; Saparov et al., 2023; Shi et al., 2023). For example, CoT heavily depends on the semantic structure of the problem to perform reasoning (Tang et al., 2023), struggles with planning and unification in deductive reasoning (Saparov and He, 2023), has difficulty generalizing to longer reasoning paths (Saparov et al., 2023), and can be easily misled by irrelevant information in the context (Shi et al., 2023). However, these observations are derived from specific reasoning tasks and prompt settings, limiting their applicability to other scenarios. In contrast, we introduce a general-purpose tool that allows users to analyze reasoning in their contexts.
Tools for analyzing chain-of-thought. To the best of our knowledge, the only existing tool for analyzing CoT is gradient-based feature attribution (Wu et al., 2023), which computes a saliency score for each input token based on the model’s output. However, these token-level saliency scores do not directly capture the thought-level, multi-step reasoning process of LLMs. Consequently, the main finding in (Wu et al., 2023) is that CoT stabilizes saliency scores on semantically relevant tokens compared to direct prompting.
Metrics designed to quantify CoT performance (Chen et al., 2024; Ton et al., 2024) can also be used to analyze the reasoning behaviors of LLMs. For instance, Ton et al. (2024) employs information gain to identify failure modes in reasoning paths, aligning with Observation 3.2 in this paper. However, our 2-D visualization offers significantly deeper insights than a single information gain metric. Additionally, the verifier derived from our tool is conceptually related to outcome-supervised reward models (Cobbe et al., 2021).

### v3

@@ 2.3 Quantitative Visualization with Metrics
Besides the qualitative visualization, we introduce three quantitative metrics to help understand the LLMs’ behavior. These metrics are defined based on the intermediate states in Sec. 2.2.
@@
Consistency. To understand whether the LLM knows the answer before generating all thoughts, we compute the consistency of state $s_{i}$ by checking whether $\bm{f}_{i}$ and $\bm{f}_{n}$ agree
$\text{Consistency}(s_{i})=\text{1}(\operatorname*{arg\,min}\bm{f}_{i}=\operatorname*{arg\,min}\bm{f}_{n}).$ (5)
@@
Uncertainty. To know how confident the LLM is about its predictions at intermediate steps, we compute the uncertainty of state $s_{i}$ as the entropy of $\bm{f}_{i}$ (note $\sum_{d\in\bm{f}_{i}}d=1$)
$\text{Uncertainty}(s_{i})=-\sum_{d\in\bm{f}_{i}}d\cdot\log d.$ (6)
@@
$\text{Perplexity}(t_{i})=p_{\text{LLM}}(t_{i}|s_{i-1})^{-1/|t_{i}|}.$ (7)
Remark 2.1.
Note that in previous works, these metrics are mainly used to evaluate the performance of language modeling on each token. We repurpose them to analyze intermediate thoughts in the trajectories, which is a new lesson for the community. Appendix B introduces related works in detail. The following section demonstrates that the LoT, containing the qualitative landscape and the quantitative metrics, is effective for automatic and scalable visualization of reasoning trajectories.
@@ 3 Results and Observations
In this section, we utilize the landscape of thoughts to analyze the reasoning behavior of LLMs by comparing the visualizations across three dimensions: (1) diverse scales and types of language models in Sec. 3.1, (2) different reasoning tasks in Sec. 3.2, and (3) various reasoning methods in Sec. 3.3. Unless stated otherwise, we employ Llama-3.1-70B with CoT as the default configuration in evaluations. All the visualizations are built upon the model’s estimation of their own thoughts.  [^5] Appendix D.1 validates each qualitative observation from LoT. Full visualizations are in Appendix E.
3.1 Comparison across Language Models
(a) Llama-3.2-1B
[…]
(b) Llama-3.2-3B
[…]
(c) Llama-3.1-8B
[…]
(d) Llama-3.1-70B
[…]
Figure 2: Comparing the LoT of different language models (with CoT on the AQuA dataset). Darker regions represent higher state density, with Refer to caption indicating incorrect answers and Refer to caption marking the correct ones. Through the reasoning trajectories, spanning from early (0-20% states) to the later stages (80-100% states), the visualization shows correct cases (bottom rows in blue) with incorrect cases (top rows in red). Metrics are calculated w.r.t. each bin, e.g., 20% - 40% of states. The reasoning accuracy of the four subfigures is: (a) 15.8%, (b) 42.0%, (c) 53.2%, and (d) 84.4%.
We study several LLMs’ behavior across parameter scales (from 1B, 3B to 70B). We run each model with CoT prompting on 50 randomly selected problems from the mathematical reasoning dataset AQuA. Their landscapes are shown in Fig. 2, from which we have the following observations.
Figure 3: The LoT of the reasoning model QwQ-32B (using CoT prompting on the AQuA dataset).
Observation 3.1 (The landscape converges faster as the model size increase).
As model parameters scale from 1B to 70B, the corresponding landscape demonstrates faster convergence to the correct answers with higher density in the last 20% states, aligning with the increasing accuracy. With more parameters, larger models can store broader knowledge (Allen-Zhu & Li, 2024). This leads to more confident solutions, demonstrated by more focused answer patterns and lower uncertainty.
Observation 3.2 (Larger models have higher consistency, lower uncertainty, and lower perplexity).
As the model size increases, the consistency increases; at the same time, the uncertainty and perplexity decrease significantly. This also aligns with the higher accuracy for the large models. [^6] Appendix D.3 presents additional analyses of the consistency metric: the consistency does not relate to the length of the trajectory. In addition, Appendix D.5 supports the validity of comparing perplexity across models.
In addition, we apply LoT to up-to-date reasoning models QwQ 32B (Team, 2025) and observe:
Observation 3.3 (Reasoning models present more-complex reasoning behaviors in landscapes.).
As shown in Fig. 3, the landscapes can capture complex reasoning patterns such as self-evaluation and self-correction. Specifically, correct trajectories tend to include more instances of self-evaluation and self-correction compared to incorrect ones. These behaviors often occur early in the reasoning process, especially when the model is far from the correct one. Compared to non-reasoning models, correct trajectories here show greater diversity, with green and yellow points more widely scattered.
@@ 3.2 Comparison across Reasoning Tasks
(a) AQuA Refer to caption
[…]
(b) MMLU Refer to caption
[…]
(c) StrategyQA Refer to caption
[…]
(d) CommonSenseQA
[…]
Figure 4: Comparing the LoT of different datasets (using Llama-3.1-70B with CoT). The accuracy of reasoning for the four subfigures is: (a) 84.4%, (b) 80.2%, (c) 75.8%, and (d) 64.8%.
Besides AQuA dataset, we include MMLU, CommonsenseQA, and StrategyQA datasets. We run the default model with CoT on 50 problems per dataset. These observations are derived from Fig. 4:
@@
The landscapes of AQuA, MMLU, and StrategyQA in Fig. 4 exhibit organized search behavior with higher state diversity, while CommonSenseQA presents concentrated search regions, reflecting direct retrieval of commonsense knowledge rather than step-by-step reasoning processes. These distinct landscape patterns demonstrate the potential to reveal underlying domain relationships across different reasoning tasks.
@@
The histograms in Fig. 4 show that the perplexity consistently increases as reasoning progresses across all datasets. Specifically, different datasets, e.g., AQuA and MMLU, show distinctly higher levels of uncertainty. As for StrategyQA, correct trajectories show increasing consistency that surpasses incorrect trajectories at around 60% states, while incorrect trajectories show decreasing consistency. However, when the trajectory is longer than the ground truth trajectory, the later stages (60-100% of states) exhibit both increasing perplexity and decreasing uncertainty.  [^7] We show detailed analysis for trajectories in StrategyQA in Appendix D.4.
3.3 Comparison across Reasoning Methods
(a) CoT Refer to caption
[…]
(b) LtM Refer to caption
[…]
(c) MCTS Refer to caption
[…]
(d) ToT Refer to caption
[…]
Figure 5: Comparing the LoT of four reasoning methods (using Llama-3.1-70B on the AQuA dataset). The reasoning accuracy is: (a) 84.4%, (b) 82.2%, (c) 75.8%, and (d) 81.6%, respectively.
Figure 6: Case Study of LoT, with Llama-3.1-8B using CoT on AQuA.
Setup. We evaluate the default model with four reasoning methods: chain-of-thought (CoT) (Wei et al., 2022), least-to-most (LtM) (Zhou et al., 2023), MCTS (Zhang et al., 2024), and tree-of-thought (ToT) (Yao et al., 2023a). We run these methods on 50 problems from AQuA and observe that:
Observation 3.6 (Cross-method comparison: Among correct reasoning trajectories, methods with faster convergence to correct answers achieve higher accuracy.).
From Fig. 5, we observe that the states scatter dispersedly at early stages and gradually converge to correct (or incorrect) answers in later stages. Here, converge means the trend of a reasoning trajectory approaching one answer. Generally, methods with more scattered landscapes (that converge more slowly) present lower accuracy than those that converge faster. For example, the blue landscape in Fig. 5(a) converges faster than the blue landscapes in Fig. 5(c), and the former is with a higher accuracy than the latter.
Observation 3.7 (Within-method comparison: For any single method, incorrect trajectories converge faster to wrong answers than correct trajectories converge to right answers.).
As can be seen from Fig. 4, failure trajectories usually converge to the wrong answers at earlier states of reasoning, e.g., 20-40% states in Fig. 4(c). By contrast, the states in the success trajectories converge to the correct answers at later 80-100% states. This implies that early states of the reasoning process can lead to any potential answers (from a model perspective), while the correct answers are usually determined at the end of reasoning trajectories. In addition, Fig. 6 showcases the corresponding text of thoughts.  [^8] In Appendix D.2, only a few incorrect trajectories (1.8%) are close to the correct answer in middle thoughts.
Observation 3.8 (Compared to failure trajectories, the intermediate states in correct trajectories have higher consistency w.r.t. the final state).
By comparing the consistency plots in Fig. 5, we found that the model generally has low consistency between the intermediate states and the final state. Notably, the consistency of wrong trajectories is significantly lower than that of correct trajectories. This implies that the reasoning process can be quite unstable. Even though decoding methods like CoT and LtM are designed to solve a problem directly (without exploration), the generated thoughts by these methods do not consistently guide the reasoning trajectory to the answer.
@@ 4 Adapting Visualization to Predictive Models
One advantage of our method is that it can be adapted to a model to predict any property users observe. Here, we show how to convert our method to a lightweight verifier for voting trajectories, following the observations in Sec. 3. Note that this methodology is not limited to verifiers. Users can use this technique to adapt the visualization tool to monitor the properties in their scenarios.
@@ 4.1 A Lightweight Verifier
Observation 3.7 and 3.8 show that the convergence speed and consistency of intermediate states can distinguish correct and wrong trajectories. Inspired by these observations, we build a model $g:{\mathbb{R}}^{(k+1)\times n}\to\{0,1\}$ to predict the correctness of a trajectory based on the state features $\{\bm{f}_{i}\}_{i=1}^{n}$ and consistency metric $\{\text{Consistency}(\bm{f}_{i})\}_{i=1}^{n}$. The insight is that the state features, used to compute the 2-D visualization, encode rich location information of the states and can be used to estimate the convergence speed. Due to the small dimensionality of these features, we parameterize $f$ with a random forest (Breiman, 2001) to avoid overfitting. We use this model as a verifier to enhance LLM reasoning (Cobbe et al., 2021). Unlike popular verifiers (Lightman et al., 2023) that involve a moderately sized language model on textual thoughts, our verifier operates on state features and is quite lightweight. We train a verifier on thoughts sampled on the training split of each dataset and apply it to vote trajectories at test time. Given $q$ trajectories sampled by a decoding method, the final prediction is produced by a weighted majority voting:
$\hat{y}$= $\operatorname*{arg\,max}_{c\in{\mathcal{C}}}\sum_{i=1}^{q}$ $\text{1}(\hat{y}^{(i)}=c)\cdot g(\{\bm{f}_{i}\}_{i=1}^{n},\{\text{Consistency}(\bm{s}_{i})\}_{i=1}^{n}).$ (8)
Figure 7: The accuracy of reasoning under different decoding methods and model scales (averaging across all four datasets). Results for each dataset are in Appendix E.
Figure 8: Demonstration of the inference-time scaling effect of the verifier. We show the voting accuracy (%) on StrategyQA scales with the number of trajectories.
(a) Transfer across datasets
(b) Transfer across models
Figure 9: Absolute accuracy changes ($\Delta$ Acc) with the verifier, compared to performance in Fig. 7 (without the verifier). The verifier is trained on each column (dataset or model) and evaluated on all rows (other datasets or models). Positive values indicate improvement in accuracy with the verifier.
@@ 4.2 Experimental Results
We evaluate our numerical verifier against an unweighted voting baseline (Wang et al., 2023b) with various models, decoding methods, and reasoning datasets. We report the accuracy here instead of commonly seen pass@k, which will be easily hacked by a simple random guess or a traverse of all candidates to obtain a high score. Detailed settings of experiments are in Appendix C. We also provide ablation studies on training the verifier and discuss and compare the variance of the verifier in Appendix D.6, and experiment on the scaling effect with different features in Appendix D.7.
@@
Effectiveness of the verifier. We first compare our verifier against the unweighted voting baseline, each applied to 10 trajectories. As shown in Fig. 7, our verifier consistently enhances the reasoning performance of all models and decoding methods, even though our verifier does not use any pre-trained language model. Notably, smaller language models (1B and 3B) show significant performance gains with the verifier’s assistance, achieving substantial improvements over their original capabilities of reasoning. We also compare the verifiers between reward-guided methods.
@@
Test-time scaling. While the improvement of the verifier seems marginal with 10 trajectories, our verifier can provide a substantial performance gain with more trajectories. We adjust the number of trajectories from 1 to 50, and plot the results of the verifier and the unweighted voting baseline in Fig. 9. Models with our verifier exhibit significantly stronger scaling behaviors, achieving over 65% accuracy. In contrast, the performance of the baseline saturated around 30% accuracy. These results suggest that our state features, which are used in both the visualization tool and the verifier, capture important information about the reasoning behavior of LLMs. Thus, the verifier can boost test-time scaling, especially in solving complex problems.
@@
Cross-dataset and cross-model transferability. One interesting property of the state features and metrics is that their shape and range are agnostic to the model and dataset, suggesting that we may deploy the verifier trained on one dataset or model in another setting. As illustrated in Fig. 9, we evaluate how the verifier transfers across reasoning datasets (e.g., train on AQuA and test on MMLU) and model scales (e.g., train on 1B model and test on 70B model). We observe some positive transfers across datasets and models. For example, a verifier trained on AQuA can improve the performance of StrategyQA by 4.5%. A verifier trained on the 70B model also improves the performance of the 3B model by 5.5%. However, some cases do not benefit from the transferring verifiers. We leave improving the transferability of the state features and metrics as future work.